OpenStudy (anonymous):

Q = mc delta t? and an example i did sample weight=.7 grams degree of water(celcius) before burning= 22 degrees degrees of water after burning= 23 degrees

OpenStudy (anonymous):

please answer

OpenStudy (nebi):

what do you want to know? \[-\xi \leftarrow \sqrt{\cos^{-1} \left[\begin{matrix}? & ? \\ ? & ?\end{matrix}\right]}\] p.s. I just wanted to test out the equation tab. It has nothing to do with the question

OpenStudy (anonymous):

umm, i just need to know if im doing it right. i did the equations, and im pretty sure a single peice of dog food does not have like, 200 Calories

OpenStudy (anonymous):

\[Q=MC \Delta T\]

OpenStudy (nebi):

Q = m*c*(delta T) Q = heat energy m = mass c = specific heat delta t = change in temperature (= final - initial temp) Remember, in a closed system, heat gained by one thing equals the heat lost by another.

OpenStudy (anonymous):

its all multiplied right?

OpenStudy (nebi):

yea. as far as i know, you did it right, the equation is correct.

OpenStudy (anonymous):

....s single peice of puppy fod has like 354 Calories!?!?!?!?

OpenStudy (nebi):

What is the change in the temperature